WAEC Mathematics 2016 Question 42

Question 42

Probability

A bag contains 5 red and 4 blue identical balls. If two balls are selected at random from the bag, one after the other, with replacement, find the probability that the first is red and the second blue.

  1. A.2/9
  2. B.5/18
  3. C.20/81
  4. D.5/9
Show the answer and explanation

Answer: C (20/81)

Explanation

With replacement: 5/9 × 4/9 = 20/81.